# A posteriori error estimation for an interior penalty virtual element method of Kirchhoff plates

### v2

@@ 7 Conclusion
Fang Feng and Yue Yu collaborated closely to shape the conceptualization, methodology, and writing of this research. Additionally, Yue Yu took charge of implementing the discrete method in the study.
@@
Data sharing is not applicable to this article as no data was generated or analyzed.
@@ Acknowledgements
Fang Feng was supported by the National Science Foundation for Young Scientists of China (No. 12401528) and also partially supported by Central Government Special Fund for Basic Scientific Research Business Expenses of Colleges and Universities, No. 30924010837. Yue Yu was partially supported by the National Science Foundation for Young Scientists of China (No. 12301561).

### v4

@@ 6.5   Formulation based on $H^{2}$-elliptic projector
All authors worked closely together on the conceptualization, methodology, and writing of this research. Additionally, Yue Yu and Yuming Hu were responsible for implementing the discrete method used in the study.
@@
No data was used for the research described in the article.
@@ Acknowledgements
Fang Feng was partially supported by the National Natural Science Foundation of China (NSFC) grant 12401528 and the Fundamental Research Funds for the Central Universities, No. 30924010837. Yue Yu was partially supported by the National Science Foundation for Young Scientists of China (No. 12301561) and the Key Project of Scientific Research Project of Hunan Provincial Department of Education (No. 24A0100). Jikun Zhao was partially supported by National Natural Science Foundation of China (No. 12371411).